# Uncovering Safety Risks of Large Language Models through Concept Activation Vector

### v3

@@ 3.2 Embedding-Level Attack Results
Overall performance. Table 1 compares our proposed SCAV with embedding-level attack baselines JRE and RepE in terms of automatically evaluated criteria, and Table 2 shows the human evaluation results. The results show that our method consistently performs the best on both datasets and LLMs, decreasing language flaws by 16% to 24%, and successfully induces the well-aligned Llama models to answer over 90% malicious instructions with useful information. We have included example cases of LLM responses in Appendix G to further illustrate the effectiveness of our method.
@@
Table 1: Automatic evaluation of embedding-level attack performance. All criteria except for ASR-keyword are evaluated by using GPT-4. The best results are in bold and the second best are underlined.
Models Methods On Advbench (%) On StrongREJECT (%)
A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$ A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$
LLaMA-2 (7B-Chat) JRE 80 76 68 70 90 72 70 70
RepE 70 90 86 44 94 98 92 24
SCAV 100 96 92 2 100 98 96 10
$\Delta$ +20 +6 +6 -42 +4 0 +4 -14
LLaMA-2 (13B-Chat) JRE 84 68 68 36 94 78 70 44
RepE 86 88 84 20 92 98 94 18
SCAV 100 98 96 0 100 100 98 2
$\Delta$ +14 +10 +12 -20 +6 +2 +4 -16
  • A-k = ASR-keyword, A-a = ASR-answer, A-u = ASR-useful, L = Language flaws, $\Delta$ = SCAV -- Best baseline
Table 2: Human evaluation of embedding-level attack performance.
Models Methods On Advbench (%) On StrongREJECT (%)
A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$
LLaMA-2 (7B-Chat) JRE 66 60 64 62 42 68
RepE 88 82 36 94 82 26
SCAV 100 92 12 96 90 8
$\Delta$ +12 +10 -24 +2 +8 -18
  • A-k = ASR-keyword, A-a = ASR-answer, A-u = ASR-useful, L = Language flaws, $\Delta$ = SCAV -- Best baseline
Impact of training data size. In this experiment, we mainly study how much training data is required for embedding-level attacks to achieve consistently high ASR-keyword. For each training data size, we randomly sample 5 subsets of data and report the average results.
As shown in Figure 3, our method only requires 5 pairs of malicious and safe instructions to achieve an average ASR-keyword that is close to 100%. Besides, the variance of our method is much smaller, indicating its stability. In comparison, the ASR-keyword of RepE is 0 when the training dataset size is 1, and both baselines perform much worse than ours at varying training data sizes due to their potentially misleading heuristics.
@@
Ablation study and sensitivity analysis. We conduct additional experiments to validate the effectiveness of important components and stability of our method. The detailed results are in Appendix  E. We summarize the major conclusions as follows: 1) we demonstrate the effectiveness of our automatic hyperparameter selection by showing that it increases ASR-useful by 2%~10% and reduces language flaws by up to 20%, compared to manually selecting better hyperparameters by humans (e.g., perturbing 9~13 layers with unified $\epsilon=-1.5$); 2) we illustrate the effectiveness of our perturbation direction by showing that our method consistently achieves better ASR-keyword compared with the baselines under varying perturbation magnitude and layers.
@@ 3.3 Prompt-Level Attack Results
Table 3: Evaluation of prompt-level attack performance.
Models Methods On Advbench (%) On StrongREJECT (%)
A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$ A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$
LLaMA-2 (7B-Chat) DeepInception 42 28 10 60 46 22 8 76
AutoDAN 24 22 14 60 30 26 10 62
SCAV 54 60 44 52 60 46 40 44
$\Delta$ +12 +32 +30 -8 +14 +20 +30 -18
LLaMA-2 (13B-Chat) DeepInception 16 8 4 58 18 16 12 54
AutoDAN 30 18 14 58 18 20 16 56
SCAV 72 46 28 58 54 48 46 42
$\Delta$ +42 +28 +14 0 +36 +28 +30 -12
  • A-k = ASR-keyword, A-a = ASR-answer, A-u = ASR-useful, L = Language flaws, $\Delta$ = SCAV -- Best baseline
Table 4: Attack transferability study: applying attack prompts learned for LLaMA to GPT-4.
Source Models Methods On Advbench (%) On StrongREJECT (%)
A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$ A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$
LLaMA-2 (7B-Chat) AutoDAN 36 28 26 68 32 22 18 82
SCAV 70 66 52 68 30 20 20 72
$\Delta$ +34 +38 +26 0 -2 -2 +2 -10
LLaMA-2 (13B-Chat) AutoDAN 34 20 24 80 12 18 16 84
SCAV 82 48 60 54 40 26 22 72
$\Delta$ +48 +28 +36 -26 +28 +8 +6 -12
  • A-k = ASR-keyword, A-a = ASR-answer, A-u = ASR-useful, L = Language flaws, $\Delta$ = SCAV -- Best baseline
Tranferability to GPT-4. Table 4 shows the results of applying prompts learned from LLaMA models to GPT-4. Our method usually performs better, improving ASR-related criteria by at most 48%, and reducing language flaws by at most 26%. This demonstrates our attack prompts learned by studying the inner workings of certain white-box models may still be useful for other black-box APIs. The potential transferability of attack prompts is also observed by previous research (Liu et al., 2023).
@@ 4 Understanding Safety Risks and Mechanisms of LLMs
The goal of this section is to provide insights into the severity of LLM safety risks and to better understand the safety mechanism of LLMs by applying our method.
@@ 4.1 Are Aligned LLMs Really Safe?
White-box LLMs. Table 5 shows the results when using SCAV to attack 7 well-known open-source LLMs (Jiang et al., 2023; Bai et al., 2023; Zheng et al., 2023; Xu et al., 2023). We can see that all LLMs provide relevant answers to more than 85% malicious instructions (ASR-answer), except for one on Advbench, which answers 78% malicious instructions. The response quality is also high, with an average ASR-useful of 87% and on average 12% language flaws. Moreover, ASR-keyword is close to 100% in most cases. This is very dangerous because 1) the performance of recently released open-source LLMs is gradually improving, and 2) almost no cost is required to obtain a response to any malicious instruction, as we do not require LLMs to be fine-tuned or large training data. This warns us that the existing alignment of the open-source LLMs can be extensively reversed, and there is an urgent need to develop effective methods to defend against current attack methods or stop open-sourcing high-performance LLMs.
Table 5: Attacking 7 well-known open-source LLMs by using SCAV. All LLMs provide relevant answers to more than 85% malicious instructions (A-a), except for one on Advbench (A-a is 78%).
Models On Advbench (%) On StrongREJECT (%)
A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$ A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$
LLaMA-2-7B-Chat 100 96 92 2 98 98 96 10
LLaMA-2-13B-Chat 100 98 96 0 100 100 98 2
LLaMA-3-8B-Instruct 100 90 82 14 100 94 92 8
Mistral-7B 100 90 84 20 94 96 92 20
Qwen-1.5-7B-Chat 100 78 66 26 100 86 78 20
Vicuna-v1.5-7B 98 94 80 12 98 86 84 22
WizardLM-2 100 96 90 8 100 90 88 10
Average 99.71 91.71 84.29 11.71 98.57 92.86 89.71 13.14
  • A-k = ASR-keyword, A-a = ASR-answer, A-u = ASR-useful, L = Language flaws
Methods On Advbench (%) On StrongREJECT (%)
A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$ A-k $\uparrow$ A-a $\uparrow$ A-u $\uparrow$ L $\downarrow$
SCAV-LLaMA-13B 82 66 60 54 40 26 22 72
SCAV-Both 96 78 80 42 52 30 36 58
All 96 84 84 28 86 54 54 44
  • A-k = ASR-keyword, A-a = ASR-answer, A-u = ASR-useful, L = Language flaws
@@ 4.2 Are Existing Unlearn Methods Really Effective?
We then study whether the existing defense methods that help LLMs unlearn harmful knowledge are effective. This is achieved by applying existing attack methods on a version of LLaMA-2-7B-Chat that has been fine-tuned to unlearn harmful knowledge by using an existing unlearn method Eraser (Lu et al., 2024). Table 7 shows that SCAV can still induce the LLM to produce many harmful responses, indicating that the unlearn method may not have fully erased harmful knowledge from the LLM, although it appears to be effective without the attack.
Table 7: After unlearning harmful knowledge by using Eraser (Lu et al., 2024), SCAV can still induce the LLM to produce many harmful responses, indicating that the unlearn method may not have fully erased harmful knowledge from the LLM, even though it appears to be effective without our attack.
@@
ASR-k (%) Harmfulness ASR-k (%) Harmfulness
@@
  • ASR-k = ASR-keyword, Harmfulness (Qi et al., 2023) is a quality criterion with a maximum score of 5
First, there may be a close relation between linear separability and the safety mechanism of LLMs. Our previous experiments have shown that 1) aligned LLMs can linearly separate embeddings from malicious and safe instructions at later layers (Figure 1), and that 2) attacks guided by the linear classifier are of high success ratio, indicating that the safety mechanism of LLMs may be well modeled by linear separability. To better understand their relation, we further attack LLaMA-2-7B-Chat on the 0th, 10th, 20th, and 30th layers. As shown in Figure 4(a), attacks on a linearly separable layer (10, 20, 30) consistently lead to an increase in ASR-keyword, while attacks on the other layer (0) do not improve ASR-keyword. Based on the results, we speculate that for every single layer, linear separability may not only indicate that LLMs understand the safety concept, but may also mean that the LLMs will use this safety concept in subsequent layers for generating responses.
@@ 4.3 How Do Aligned LLMs Differentiate Malicious Instructions from Others?
Figure 4: Unveiling the safety mechanism of LLMs by (a) attacking a single layer; (b) attacking multiple layers, and (c) transferring embedding-level attacks to other white-box LLMs.
Second, different layers may have modeled the safety mechanism from related but different perspectives. Figure 4(b) shows the value of $P_{\text{m}}$ when attacking different layers of LLaMA-2-7B-Chat. We have two observations. (1) While attacking a single layer (Layer 10) results in a low $P_{\text{m}}$ at the current layer, $P_{\text{m}}$ subsequently increases on the following layers. This means that later layers somehow gradually correct the attack by leveraging existing information of the embedding, potentially because it models the safety mechanism from a different perspective. (2) We observe that when more layers are perturbed (e.g., layers 10-13), $P_{\text{m}}$ at later layers can no longer be corrected by the LLM. This indicates that a limited number of layers may jointly determine the overall safety mechanism from different perspectives.
Finally, different white-box LLMs may share some commonalities in their safety mechanism. Figure 4(c) showcases ASR-keyword when applying embedding-level attacks from one white-box model to another. We can see that the ASR-keyword is sometimes quite large. This indicates that the safety mechanism of LLMs may have certain commonalities and that SCAV may have characterized this commonality in some sense. However, there is still a lack of clear understanding of when it can transfer and why.
@@ 5 Conclusion
Limitation. Although our method performs well at both embedding and prompt levels, we lack an in-depth exploration of the transferability mechanism of perturbation vectors and attack prompts. We believe this is a potential future direction toward the construction of responsible AI.
@@
Appendix A Evaluation Rules
A.1 ASR-keyword
ASR-keyword uses a keyword-matching algorithm to determine whether the attack is successful. Empirically, LLM usually responds in a fixed pattern when refuses to follow the user’s instructions (Zou et al., 2023b). Therefore, once any of the refusal keywords we have selected (see Table 8) appears in LLM’s response, the corresponding test case of this response will be considered failed, and vice versa, it will be considered successful.

### v5

@@ 3.2 Embedding-Level Attack Results
Overall performance. Table 1 compares our proposed SCAV with embedding-level attack baselines JRE and RepE in terms of automatically evaluated criteria, and Table 2 shows the human evaluation results. The results show that our method consistently performs the best on both datasets and LLMs, decreasing language flaws by 16% to 24%, and successfully induces the well-aligned Llama models to answer over 90% malicious instructions with useful information. We have included example cases of LLM responses in Appendix I to further illustrate the effectiveness of our method.
@@
Table 1: Automatic evaluation of embedding-level attack performance. All criteria except for ASR-keyword are evaluated by GPT-4. The best results are in bold and the second best are underlined. $\Delta$ = SCAV -- Best baseline.
Models Methods Results on (Advbench / StrongREJECT), %
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language flaws $\downarrow$
LLaMA-2 (7B-Chat) JRE 80 / 90 76 / 72 68 / 70 70 / 70
RepE 70 / 94 90 / 98 86 / 92 44 / 24
Soft prompt 56 / 64 50 / 44 40 / 38 62 / 66
SCAV 100 / 100 96 / 98 92 / 96 2 / 10
$\Delta$ +20 / +4 +6 / 0 +6 / +4 -42 / -14
LLaMA-2 (13B-Chat) JRE 84 / 94 68 / 78 68 / 70 36 / 44
RepE 86 / 92 88 / 98 84 / 94 20 / 18
Soft prompt 80 / 74 66 / 28 50 / 28 44 / 68
SCAV 100 / 100 98 / 100 96 / 98 0 / 2
$\Delta$ +14 / +6 +10 / +2 +12 / +4 -20 / -16
Table 2: Human evaluation of embedding-level attack performance. $\Delta$ = SCAV -- Best baseline.
Models Methods Results on (Advbench / StrongREJECT), %
ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language flaws $\downarrow$
LLaMA-2 (7B-Chat) JRE 66 / 62 60 / 42 64 / 68
RepE 88 / 94 82 / 82 36 / 26
SCAV 100 / 96 92 / 90 12 / 8
$\Delta$ +12 / +2 +10 / +8 -24 / -18
@@
Impact of training data size. In this experiment, we mainly study how much training data is required for embedding-level attacks to achieve consistently high ASR-keyword. For each training data size, we randomly sample 5 subsets of data and report the average results. As shown in Figure 3, our method only requires 5 pairs of malicious and safe instructions to achieve an average ASR-keyword that is close to 100%. Besides, the variance of our method is much smaller, indicating its stability. In comparison, the ASR-keyword of RepE is 0 when the training dataset size is 1, and both baselines perform much worse than ours at varying training data sizes due to their potentially misleading heuristics.
Ablation study and sensitivity analysis. We conduct additional experiments to validate the effectiveness of important components and stability of our method. The detailed results are in Appendix  F. We summarize the major conclusions as follows:
  • We demonstrate the effectiveness of our automatic hyperparameter selection by showing that it increases ASR-useful by 2%~10% and reduces language flaws by up to 20%, compared to manually selecting better hyperparameters by humans (e.g., perturbing 9~13 layers with unified $\epsilon=-1.5$).
  • We illustrate the effectiveness of our perturbation direction by showing that our method consistently achieves better ASR-keyword compared with the baselines under varying perturbation magnitude and layers.
@@ 3.3 Prompt-Level Attack Results
Table 3: Evaluation of prompt-level attack performance. $\Delta$ = SCAV -- Best baseline.
Models Methods Results on (Advbench / StrongREJECT), %
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language flaws $\downarrow$
LLaMA-2 (7B-Chat) DeepInception 42 / 46 28 / 22 10 / 8 60 / 76
AutoDAN 24 / 30 22 / 26 14 / 10 60 / 62
GCG 28 / 26 32 / 26 10 / 16 76 / 72
SCAV 54 / 60 60 / 46 44 / 40 52 / 44
$\Delta$ +12 / +14 +28 / +20 +30 / +24 -8 / -18
LLaMA-2 (13B-Chat) DeepInception 16 / 18 8 / 16 4 / 12 58 / 54
AutoDAN 30 / 18 18 / 20 14 / 16 58 / 56
GCG 40 / 34 24 / 18 10 / 16 58 / 80
SCAV 72 / 54 46 / 48 28 / 46 58 / 42
$\Delta$ +32 / +20 +22 / +28 +14 / +30 0 / -12
Tranferability to GPT-4. Table 4 shows the results of applying prompts learned from LLaMA models to GPT-4. Our method usually performs better, improving ASR-related criteria by at most 48%, and reducing language flaws by at most 26%. This demonstrates our attack prompts learned by studying the inner workings of certain white-box models may still be useful for other black-box APIs. The potential transferability of attack prompts is also observed by previous research [14].
Table 4: Attack transferability study: applying attack prompts learned for LLaMA to GPT-4. $\Delta$ = SCAV -- Best baseline.
Source Models Methods Results on (Advbench / StrongREJECT), %
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language flaws $\downarrow$
LLaMA-2 (7B-Chat) AutoDAN 36 / 32 28 / 22 26 / 18 68 / 82
GCG 4 / 8 4 / 16 2 / 16 92 / 90
SCAV 70 / 30 66 / 20 52 / 20 68 / 72
$\Delta$ +34 / -2 +38 / -2 +26 / +2 0 / -10
LLaMA-2 (13B-Chat) AutoDAN 34 / 12 20 / 18 24 / 16 80 / 84
GCG 2 / 8 0 / 12 0 / 10 98 / 88
SCAV 82 / 40 48 / 26 60 / 22 54 / 72
$\Delta$ +48 / +28 +28 / +8 +36 / +6 -26 / -12
@@ 4 Understanding Safety Risks and Mechanisms of LLMs
The goal of this section is to provide insights into the severity of LLM safety risks and to better understand the safety mechanisms of LLMs by applying our method.
@@ 4.1 Are Aligned LLMs Really Safe?
White-box LLMs. Table 5 shows the results when using SCAV to attack 7 well-known open-source LLMs [36; 37; 38; 39]. We can see that all LLMs provide relevant answers to more than 85% malicious instructions (ASR-answer), except for one on Advbench, which answers 78% malicious instructions. The response quality is also high, with an average ASR-useful of 87% and on average 12% language flaws. Moreover, ASR-keyword is close to 100% in most cases. This is very dangerous because 1) the performance of recently released open-source LLMs is gradually improving, and 2) almost no cost is required to obtain a response to any malicious instruction, as we do not require LLMs to be fine-tuned or large training data. This warns us that the existing alignment of the open-source LLMs can be extensively reversed, and there is an urgent need to develop effective methods to defend against current attack methods or stop open-sourcing high-performance LLMs.
Table 5: Attacking 7 well-known open-source LLMs by using SCAV. All LLMs provide relevant answers to more than 85% malicious instructions (ASR-answer), except for one on Advbench (ASR-answer is 78%).
Models Results on (Advbench / StrongREJECT), %
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language flaws $\downarrow$
LLaMA-2-7B-Chat 100 / 98 96 / 98 92 / 96 2 / 10
LLaMA-2-13B-Chat 100 / 100 98 / 100 96 / 98 0 / 2
LLaMA-3-8B-Instruct 100 / 100 90 / 94 82 / 92 14 / 8
Mistral-7B 100 / 94 90 / 96 84 / 92 20 / 20
Qwen-1.5-7B-Chat 100 / 100 78 / 86 66 / 78 26 / 20
Vicuna-v1.5-7B 98 / 98 94 / 86 80 / 84 12 / 22
WizardLM-2 100 / 100 96 / 90 90 / 88 8 / 10
Average 99.71 / 98.57 91.71 / 92.86 84.29 / 89.71 11.71 / 13.14
@@
Methods Results on (Advbench / StrongREJECT), %
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language flaws $\downarrow$
SCAV-LLaMA-13B 82 / 40 66 / 26 60 / 22 54 / 72
SCAV-Both 96 / 52 78 / 30 80 / 36 42 / 58
All 96 / 86 84 / 54 84 / 54 28 / 44
@@ 4.2 Are Existing Unlearn Methods Really Effective?
We then study whether the existing defense methods that help LLMs unlearn harmful knowledge are effective. This is achieved by applying existing attack methods on a version of LLaMA-2-7B-Chat that has been fine-tuned to unlearn harmful knowledge by using an existing unlearn method Eraser [21]. Table 7 shows that SCAV can still induce the LLM to produce many harmful responses, indicating that the unlearn method may not have fully erased harmful knowledge from the LLM, although it appears to be effective without the attack. Furthermore, we find that existing defense methods might not effectively mitigate the proposed embedding-level attacks (see Appendix G).
Table 7: After unlearning harmful knowledge by using Eraser [21], SCAV can still induce the LLM to produce many harmful responses, indicating that the unlearn method may not have fully erased harmful knowledge from the LLM, even though it appears to be effective without our attack. Harmfulness [40] is a quality criterion with a maximum score of 5.
@@
ASR-keyword (%) Harmfulness ASR-keyword (%) Harmfulness
@@
Figure 4: Unveiling the safety mechanisms of LLMs by (a) attacking a single layer; (b) attacking multiple layers, and (c) transferring embedding-level attacks to other white-box LLMs.
First, there may be a close relation between linear separability and the safety mechanisms of LLMs. Our previous experiments have shown that 1) aligned LLMs can linearly separate embeddings from malicious and safe instructions at later layers (Figure 1), and that 2) attacks guided by the linear classifier are of high success ratio, indicating that the safety mechanisms of LLMs may be well modeled by linear separability. To better understand their relation, we further attack LLaMA-2-7B-Chat on the 0th, 10th, 20th, and 30th layers. As shown in Figure 4(a), attacks on a linearly separable layer (10, 20, 30) consistently lead to an increase in ASR-keyword, while attacks on the other layer (0) do not improve ASR-keyword. Based on the results, we speculate that for every single layer, linear separability may not only indicate that LLMs understand the safety concept, but may also mean that the LLMs will use this safety concept in subsequent layers for generating responses.
Second, different layers may have modeled the safety mechanisms from related but different perspectives. Figure 4(b) shows the value of $P_{\text{m}}$ when attacking different layers of LLaMA-2-7B-Chat. We have two observations. First, while attacking a single layer (Layer 10) results in a low $P_{\text{m}}$ at the current layer, $P_{\text{m}}$ subsequently increases on the following layers. This means that later layers somehow gradually correct the attack by leveraging existing information of the embedding, potentially because it models the safety mechanisms from a different perspective. Second, we observe that when more layers are perturbed (e.g., layers 10-13), $P_{\text{m}}$ at later layers can no longer be corrected by the LLM. This indicates that a limited number of layers may jointly determine the overall safety mechanisms from different perspectives.
Finally, different white-box LLMs may share some commonalities in their safety mechanisms. Figure 4(c) showcases ASR-keyword when applying embedding-level attacks from one white-box model to another. We can see that the ASR-keyword is sometimes quite large. This indicates that the safety mechanisms of LLMs may have certain commonalities and that SCAV may have characterized this commonality in some sense. However, there is still a lack of clear understanding of when it can transfer and why.
@@ 5 Conclusion
Limitation. Although our method performs well at both embedding and prompt levels, we lack an in-depth exploration of the transferability mechanisms of perturbation vectors and attack prompts. We believe this is a potential future direction toward the construction of responsible AI.
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC) (NO. 62476279), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, Kuaishou, and the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China No. 24XNKJ18. This work was partially done at Beijing Key Laboratory of Big Data Management and Analysis Methods and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education. This research was supported by Public Computing Cloud, Renmin University of China.
@@ Acknowledgements
Appendix A Pipeline Overview
To enhance understanding, we present a pipeline demonstration for conducting embedding-level and prompt-level attacks using SCAVs, as illustrated in Figure 5.
The pipeline consists of three conceptual LLMs (A, B, and C). In our primary experiments, we assume A, B, and C are identical (A = B = C). For embedding-level transferability settings, we assume that LLMs A and B share the same embedding dimensions. In contrast, for prompt-level transferability settings, no additional assumptions are made regarding A, B, or C.
Figure 5: A Pipeline Demonstration for Conducting Embedding-Level and Prompt-Level Attacks Using SCAVs.
Appendix B Evaluation Rules
B.1 ASR-keyword
ASR-keyword uses a keyword-matching algorithm to determine whether the attack is successful. Empirically, LLM usually responds in a fixed pattern when refuses to follow the user’s instructions [14]. Therefore, once any of the refusal keywords we have selected (see Table 8) appears in LLM’s response, the corresponding test case of this response will be considered failed, and vice versa, it will be considered successful.